### v1

# Bokeh Diffusion: Defocus Blur Control in Text-to-Image Diffusion Models

Abstract
Recent advances in large-scale text-to-image models have revolutionized creative fields by generating visually captivating outputs from textual prompts; however, while traditional photography offers precise control over camera settings to shape visual aesthetics—such as depth-of-field—current diffusion models typically rely on prompt engineering to mimic such effects. This approach often results in crude approximations and inadvertently altering the scene content. In this work, we propose Bokeh Diffusion, a scene-consistent bokeh control framework that explicitly conditions a diffusion model on a physical defocus blur parameter. By grounding depth-of-field adjustments, our method preserves the underlying scene structure as the level of blur is varied. To overcome the scarcity of paired real-world images captured under different camera settings, we introduce a hybrid training pipeline that aligns in-the-wild images with synthetic blur augmentations. Extensive experiments demonstrate that our approach not only achieves flexible, lens-like blur control but also supports applications such as real image editing via inversion.
Figure 1: Bokeh Diffusion enables precise, scene-consistent bokeh transitions in text-to-image diffusion models. Images are explicitly conditioned on a defocus parameter: the top row is generated with a low blur value to produce sharp, nearly all-in-focus outputs, while the bottom row is conditioned toward a high blur value, resulting in a strong bokeh effect.
1 Introduction
Photography has long offered a rich set of camera controls for shaping visual aesthetics. By varying the lens aperture and focal distance, photographers can decide whether a scene appears deeply in focus or softly blurred in the foreground or background. In professional workflows, these lens-based adjustments are integral to crafting the photograph’s mood and composition. Yet, in the quickly evolving domain of text-to-image generation, reproducing such camera effects is surprisingly difficult. Existing approaches typically rely on prompt engineering (e.g., adding “f/1.8 aperture” to the textual description), which often yields only a crude approximation of optical blur and unintentionally alters the scene content. As a result, changing from a “wide aperture” to a “narrow aperture” might yield not just a different depth-of-field but a new arrangement of objects.
In this work, we seek to achieve scene-consistent bokeh control in text-to-image diffusion models. A key obstacle is that real-world images with paired camera settings for the same scene (e.g., identical viewpoint at f/1.8 and f/16) are scarce. To address this, we align in-the-wild images (which contain lens-related metadata) and synthetic blur augmentations, enabling the model to learn from both structured focus–defocus pairs and authentic in-the-wild lens blur. Through this hybrid data pipeline, our model learns to better disentangle the image scene from its depth-of-field while closely mimicking lens-based defocus blur.
Beyond aligning real-world and synthetic blur augmentations, our framework conditions the diffusion process on an explicit physical bokeh parameter and employs specialized attention mechanisms to preserve scene layout. Specifically, we employ lightweight cross-attention adapters that inject blur cues into the model without overwriting textual content, enabling fine-grained adjustments in depth-of-field. Additionally, our grounded self-attention is trained with pivot-based pairs, so that altering the blur intensity in one image respects the structure of another reference image. Together, these components disentangle blur from the underlying scene, allowing us to transition between shallow and deep focus without inadvertently reshaping the layout.
Figure 2: Comparison with blur renderer BokehMe [30].
Compared to post-processing approaches that add blur to images based on depth estimation [30], our approach has several key advantages. First, it enables direct bokeh control, allowing users to achieve the desired bokeh strength immediately without post-processing. Second, given an initial image, our approach allows for adjusting the bokeh strength in both directions—either increasing or decreasing it—whereas post-processing methods struggle to remove blur. Third, leveraging the image prior of the pretrained diffusion model and the hybrid dataset, our method often produces more natural bokeh effects, especially in challenging areas such as thin structures (see Figure 2).
In summary, (1) we propose Bokeh Diffusion, a scene-consistent bokeh control framework for text-to-image diffusion, allowing physically motivated depth-of-field changes without content modification; (2) we introduce a curated dataset and hybrid training pipeline that leverage EXIF-based cues from in-the-wild data and synthetic blur augmentations, furnishing the model with contrastive defocus states; (3) we present extensive experiments to validate how our approach preserves scene composition while enabling lens-like blur adjustments, something that prompt-based methods struggle to achieve; (4) we present additional native applications, such as real image editing via inversion.
2 Related Work
Text-to-image diffusion.
Diffusion models [37, 16] have become the dominant approach in text-to-image (T2I) generation due to their high sample quality and diversity. Early works such as GLIDE [26] introduced text-conditioned diffusion with a cascaded architecture for improved resolution. Stable Diffusion [33] shifted generation to a lower-dimensional latent space, enabling efficient high-resolution synthesis. Other prominent works include [35, 3, 31, 7, 8]. More recently, Stable Diffusion 3 [10] and FLUX [19] have set a new state-of-the-art, by scaling diffusion transformer [28, 2] backbones with rectified flows [22, 21]. Despite their advancements in prompt fidelity and scene composition, these models struggle with fine-grained photographic controls, such as depth-of-field manipulation.
Other recent works have also explored enhancing controllability in diffusion models. DreamBooth [34] enables subject-driven generation, while ControlNet [51] incorporates structural conditions like pose or edge maps. T2I-Adapter [25] and IP-Adapter [49] facilitate reference image-driven generation. Further studies reveal that self- and cross-attention layers encode structural and semantic information, which has been leveraged for editing [14, 6], style adaptation [15, 42], and appearance transfer [1].
Camera-conditioned diffusion models.
Most existing works on camera control in diffusion models focus on extrinsic parameters, such as camera position and viewing angle [18, 17, 9], or trajectory in video generation [44, 13, 38, 46, 47]. In contrast, intrinsic camera control has received comparatively little attention. Curved Diffusion [39] explores conditioning T2I diffusion on diverse optical geometries to simulate lens-induced distortions, such as fisheye and panoramic effects. Additionally, some methods achieve zoom effects by implicitly modeling focal length [43, 9].
Generative Photography [50] is a concurrent work that disentangles intrinsic camera control by modeling different camera parameters across the temporal axis of a text-to-video diffusion model. In contrast, our method remains within a T2I paradigm and learns to transfer the depth-of-field from one image while anchoring the scene structure of another. This allows our approach to be directly applied to widely used T2I models and avoids the overhead of constantly generating several frames at inference time.
3 Method
3.1 Background
Diffusion models [37, 16]
are generative models that learn to synthesize data samples from Gaussian noise. In the forward process, diffusion models gradually add noise to the data $x_{0}\sim q\left(x_{0}\right)$, formalized by a Markov chain:
$q\left(x_{1:T}|x_{0}\right)=\prod\limits_{t=1}^{T}q\left(x_{t}|x_{t-1}\right),$ (1)
$q\left(x_{t}|x_{t-1}\right)=\mathcal{N}\left(x_{t};\sqrt{\alpha_{t}}x_{t-1},\beta_{t}\bm{I}\right),$ (2)
where $\beta_{t}$ is the noise schedule and $\alpha_{t}=1-\beta_{t}$.
In the reverse process, data is synthesized by a Gaussian model that learns to approximate the denoising step:
$p_{\theta}\left(x_{t-1}|x_{t}\right)=\mathcal{N}\left(x_{t-1};\mu_{\theta}\left(x_{t},t\right),\sigma_{t}^{2}\bm{I}\right).$ (3)
The mean $\mu_{\theta}\left(x_{t},t\right)$ can be derived to learn a noise prediction network $\epsilon_{\theta}$, which is often trained with MSE loss:
$\min_{\theta}\quad\mathbb{E}_{t,x_{0},\epsilon}\|\epsilon-\epsilon_{\theta}\left(x_{t},t\right)\|_{2}^{2},$ (4)
where $x_{t}=\sqrt{\overline{\alpha}_{t}}x_{0}+\sqrt{1-\overline{\alpha}_{t}}\epsilon$ and $t\sim\{1,2,\dots,T\}$.
For conditional generation, such as text-to-image (T2I), we model the distribution $q\left(x_{0}|c\right)$. The noise prediction network can be adapted by incorporating the condition $c$ during training, typically via the attention mechanism.
Attention mechanism in T2I diffusion.
Formally, the attention mechanism can be expressed as a weighted sum of values $V$, with the weights determined by the compatibility between a query $Q$ and a set of keys $K$:
$\text{Attention}(Q,K,V)=\text{Softmax}\Bigl{(}\tfrac{QK^{\top}}{\sqrt{d}}\Bigr{)}V.$ (5)
In most T2I diffusion models [33, 35, 31, 7, 8], $Q$ is projected from the spatial features, while $K$ and $V$ are projected either from the same spatial features (self-attention) or from text embeddings (cross-attention). Overall, these layers hold substantial information about the structure and content, allowing diffusion models to maintain coherent layouts while generating contextually relevant details.
Thin lens model.
As a camera captures a scene onto a 2D imaging plane, some parts may be in focus and others not. Given a circular aperture, the image of any point source not located at the focal distance is a small disk. This disk is often referred to as the circle of confusion (CoC), and its diameter $d$ depends on the scene and camera settings [20, 41]. Specifically, $d$ can be calculated by:
$d=\frac{f^{2}}{\,N\left(S_{1}-f\right)\,}\frac{\,\lvert S_{2}-S_{1}\rvert\,}{S_{2}},$ (6)
where $f$ is the focal length, $N$ is the aperture (i.e., F-number), $S_{1}$ is the focus distance, $S_{2}$ is the subject distance.
Defocus blur rendering.
In blur rendering methods [29, 48], the idea is to scatter each pixel in its neighboring areas where the distance between them is less than its blur radius. As discussed in [48, 40, 30], given the disparity of a pixel, its blur radius can be computed by:
$r=K\,\lvert\frac{1}{S_{1}}-\frac{1}{S_{2}}\rvert\,.$ (7)
Such methods are capable of rendering realistic bokeh effects in depth-continuous areas [30]. However, they tend to suffer from color-bleeding artifacts at depth discontinuities. Recent rendering methods mainly focus on improving this limitation, e.g., using neural renderers to fix erroneous areas [30], or layered occlusion-aware bokeh rendering [36].
3.2 Leveraging In-the-wild Image Augmentation
In principle, real photographs with naturally occurring lens blur provide the most faithful representation of defocus blur in varied scenes. However, major limitations arise when relying exclusively on in-the-wild (ITW) data. First, images of the same scene captured at multiple camera settings (e.g., identical viewpoint at f/1.8 and f/16) are exceedingly rare. Consequently, the model can struggle to disentangle the defocus blur from the underlying scene arrangement, leading to local minima. Second, real images often exhibit shooting preference bias (e.g., portrait photographers opting for wide apertures, landscape photographers going for narrow ones).
Dataset collection (Fig. 3).
We address these challenges by curating a corpus of raw photographs, each with (a) estimated depth maps, (b) foreground masks, (c) EXIF metadata, and (d) text captions. We filter out images lacking reliable lens info, unreasonable foreground area, small depth range, etc. Finally, we perform balancing based on our blur annotation obtained from the EXIF metadata (see supplementary). This step mitigates the strong long-tail bias in real photographs, providing a more uniform representation.
Synthetic blur augmentation (Fig. 4).
Even so, the scarcity of truly paired images with multiple bokeh states per scene remains a critical bottleneck. To overcome this, we synthesize additional defocus examples by rendering artificial blur on nearly all-in-focus images. Specifically, we estimate a per-pixel disparity map and then sample a blur parameter $K$ to generate contrastive “focus–defocus” states of the same image. Such synthetic data helps the model avoid local minima: it sees direct evidence that the same object layout can be rendered sharply or with shallow focus, thus faster learning to separate bokeh from scene semantics.
Mixed dataset.
In practice, combining real ITW images (offering natural-looking lens blur) with synthetic bokeh augmentations (providing explicit contrastive pairs) provides an effective balance. The real data anchors the network’s blur realism and provides diversity, while the synthetic pairs accelerate convergence and improve disentanglement of scene structure from camera-induced defocus. More details about aligning the bokeh annotation between ITW and rendered data in the supplementary material.
Figure 3: Data collection pipeline.
Figure 4: In-the-wild image augmentation.
Figure 5: Overview of Bokeh Diffusion’s conditioning and scene-consistency mechanisms. Bokeh Cross-Attention (Sec. 3.3), applied at deeper layers of the text-to-image diffusion model, injects a defocus conditioning signal, decoupled from semantic guidance. Grounded Self-Attention (Sec. 3.4) ensures fine-grained scene consistency across different defocus levels, where the pivot image’s queries anchor the scene structure, while shared keys allow the model to disentangle defocus blur from scene composition.
3.3 Defocus Blur Conditioning
Bokeh cross-attention.
Inspired by previous work [49], we employ decoupled cross-attention to inject our defocus blur condition into the existing T2I diffusion pipeline without disrupting scene semantics. Specifically, for each cross-attention layer in the decoder of the UNet, we add a new pair of key-value projections $W^{\prime}_{k},W^{\prime}_{v}$ tied to a lightweight bokeh feature $c_{b}$. This feature is obtained by passing the scalar defocus parameter through a small MLP. The original text-based keys and values ($K,V$) remain frozen. Meanwhile, the new bokeh cross-attention is computed with the same query $Q$ but keys and values $K_{b}=c_{b}W_{b}^{k},\ V_{b}=c_{b}W_{b}^{v}$, then linearly combined with the text cross-attention output:
$h^{new}=\underbrace{\text{Attention}(Q,K,V)}_{\text{text cross-attn}}+\lambda\cdot\underbrace{\text{Attention}(Q,K_{b},V_{b})}_{\text{bokeh cross-attn}}.$ (8)
Empirically, attaching our adapter to the deeper cross-attention layers sufficiently shapes the final texture while minimizing composition leakage. Figure 5 illustrates how text and bokeh features are combined for defocus control without disrupting scene semantics.
Localized cross-attention.
Additionally, we adapt a cross-attention localization strategy [45] to steer the bokeh cross-attention toward the background region during training. Specifically, we downsample each foreground mask and measure the average bokeh cross-attention in the foreground and background. Analogously to [45], we then add a small regularization term that encourages higher attention values in the background and lower in the foreground.
3.4 Imposing Scene-consistent Bokeh Generation
Grounded self-attention.
While bokeh cross-attention effectively controls defocus blur for a single image, we observe unwanted content shifts when applying different blur levels to the same scene. Simply using a structure-fixing approach at inference, such as cross-image attention [1], can enforce consistency but often conflates scene composition and blur, leading to a limited bokeh range. To address this, we introduce a grounded self-attention mechanism (see Figure 5). Concretely, we designate a random pivot image in each contrastive batch as a structural anchor for the other images: the pivot’s queries enforce a consistent scene layout, while the keys are concatenated from both pivot and target images to align each defocus level with the pivot’s content. Exposing the model to such paired examples teaches it to disentangle bokeh intensity from scene contents.
Formally, let $Q_{i},K_{i},V_{i}$ be the query, key, and value projections from the deep features $h_{i}$ of image $\mathcal{I}_{i}$, and let $Q_{p},K_{p},V_{p}$ be those corresponding to a pivot image $\mathcal{I}_{p}$. We replace the usual self-attention with:
$\text{Attention}\bigl{(}Q_{p},\,[K_{i},\,K_{p}],\,[V_{i},\,V_{i}]\bigr{)},$ (9)
[…]
To enable this cross-image guidance, we match the noise and timesteps for $\mathcal{I}_{i}$ and $\mathcal{I}_{p}$ during training, so both images undergo synchronized diffusion. Consequently, the backward (denoising) process for $\mathcal{I}_{i}$ can be expressed as:
$z_{t-1}\;=\;\hat{\epsilon}_{\theta}\bigl{(}z_{t},\,c,\,c_{b},\,t;\,\{Q_{t}^{p},K_{t}^{p}\}\bigr{)},$ (10)
where $z_{t-1}$ denotes the predicted latent at timestep $(t-1)$, $c$ is the textual context, $c_{b}$ the bokeh condition, and $\{Q_{t}^{p},K_{t}^{p}\}$ are the pivot’s queries and keys at timestep $t$.
Color transfer.
Although our model mitigates most inconsistencies, we observe occasional biases in lighting or color between different defocus levels. To address this, we optionally apply a post-processing step that aligns the color distribution of the target bokeh image with the pivot image.
Inspired by the classic color transfer method of Reinhard et al. [32], we operate in the $L^{*}a^{*}b^{*}$ color space, leveraging the fact that blur primarily alters image variance rather than its mean. Let $\mathcal{I}_{p}$ be the pivot (source) image and $\mathcal{I}_{i}$ the target image. We update each channel $c$ of $\mathcal{I}_{i}$ as:
$\mathcal{I}_{i}^{\prime,c}\;=\;\tfrac{\sigma_{p}^{c}}{\sigma_{i}^{c}}\,(\,\mathcal{I}_{i}^{c}-\mu_{i}^{c})\;+\;\mu_{p}^{c},$ (11)
thus matching channel-wise color statistics. This preserves the mean while scaling the variance, harmonizing the color in the image without degrading bokeh (see supplementary).
4 Experiments
Figure 6: Visual comparison with prominent text-to-image models SD1.5 [33], SD3.5 [10] and FLUX [19]. Attempts to guide these baselines toward a target bokeh level often yield crude approximations of defocus and may inadvertently modify the scene arrangement. In contrast, our method achieves both accurate blur control and scene consistency. Best viewed zoomed in.
4.1 Implementation Details
Dataset collection.
We build on the Megalith-10M dataset [5], consisting of roughly 10 million links to Flickr [^1] https://www.flickr.com/ images tagged as “photos” and licensed under public-domain-friendly terms. Flickr’s rich EXIF metadata makes it a valuable resource for research into lens-based blur effects. For our synthetic blur augmentations, we employ BokehMe [30] to generate defocus blur on a selected subset of nearly all-in-focus images from our curated dataset. Foreground masks are produced via briaai/RMBG-2.0 [53], and metric depth estimates come from Depth Pro [4]. The resulting dataset has 15K samples, of which around 10% are nearly all-in-focus.
Training and inference settings.
We train on a mixed corpus of in-the-wild and synthetically augmented data. Each training step samples either a real batch or a synthetic batch with equal probability. In contrastive batches, we perform grounded generation, while in-the-wild batches use conventional self-attention. We use the AdamW optimizer [23] with learning rate $1e^{-4}$ and train for approximately three hours on four RTX 4090 GPUs. Following [11], we apply a noise offset of $0.04$. We also set a min-SNR of $2.0$ [12]. At inference, we usually run $50$ denoising steps with DPM-Solver [24]; of these, about 30 steps employ grounded self-attention. Inference results for our method and SD1.5 are based on community checkpoints [^2] SG161222/Realistic_Vision_V5.1333cyberdelia/CyberRealistic.
4.2 Scene-consistent Defocus Blur Control
Figure 7: Editing depth-of-field with Bokeh Diffusion. (a) Generated image editing: By selecting a pivot image with a given bokeh level, the model can flexibly increase or decrease defocus while maintaining scene composition. (b) Real image editing: A real photograph is inverted and denoised with a new bokeh condition, preserving structural details while modifying depth-of-field.
We compare our method against three text-to-image (T2I) diffusion models: FLUX [19], SD3.5 [10], and SD1.5 [33] (our base model). To introduce bokeh conditions in the baselines, we add short descriptors to the text prompt. As a reference, we also include blur renderings from BokehMe [30]. We benchmark on 20 different prompts, each with 5 different bokeh conditions, totaling 100 samples. The evaluation metrics focus on two aspects:
  • Accuracy: We measure how each method’s bokeh aligns with the target blur parameter. Specifically, we compute the average Laplacian variance trend for images in a scene under different blur conditions, then compute Pearson correlation with the reference from BokehMe (LVCorr). A higher correlation indicates more faithful defocus control. Additionally, we employ GPT-4o [27] to estimate the bokeh level of a batch of images and calculate the Pearson correlation with its conditioning sequence.
  • Consistency: We examine whether the blur level causes unintended shifts in the scene. We use LPIPS [52] to quantify how distant images with same prompt but different blur conditions are. It is important to note that lower is not always better, since the change in depth-of-field should affect perceptual distance to some extent. We consider best the closest result to the reference. Additionally, we employ GPT-4o [27] to rate how consistent a batch of images remain despite bokeh variations.
Table 1: Quantitative comparison against baselines. Accuracy is measured by blur-trend correlation to reference (LVCorr) and GPT-4o estimations. Consistency by LPIPS (closer to reference is better) and GPT-4o scores. Best in bold, second-best underlined.
Method Accuracy Consistency
LVCorr\,$\uparrow$ GPT-4o\,$\uparrow$ LPIPS GPT-4o\,$\uparrow$
Reference 1.000 0.941 0.075 96.65
SD3.5 -0.088 0.153 0.161 80.5
FLUX -0.041 0.336 0.177 78.25
SD1.5 -0.410 0.362 0.314 68.75
Ours 0.901 0.960 0.080 94.75
Table 1 shows that our approach outperforms all baselines by a significant margin in both accuracy and consistency. Standard T2I diffusion models struggle to achieve reliable control over depth-of-field, as reflected in their low correlation scores for blur trends. Furthermore, modifying the condition frequently alters the scene, as evidenced by the consistency scores. Figure 6 offers a qualitative comparison: while the vanilla T2I models often alter the subject arrangement or fail to match the intended blur strength, our method faithfully changes bokeh levels while preserving the scene. Finally, image–text alignment measured via CLIP embeddings yields comparable scores (around 0.330) for our method and the base SD1.5 model, indicating that our additional modules do not compromise the generative quality of the underlying model.
Table 2: Quantitative ablation study.
Method Dataset Accuracy Consistency
ITW SYN LVCorr GPT-4o LPIPS GPT-4o
Reference 1.000 0.941 0.075 96.65
Bokeh Cross-Attn 0.851 0.304 0.098 90.50
Bokeh Cross-Attn 0.887 0.618 0.202 80.00
Bokeh Cross-Attn 0.867 0.690 0.171 80.00
w/ Cross-Img Attn 0.871 0.759 0.034 96.65
Ours 0.901 0.960 0.080 94.75
Generated image editing.
Our framework accommodates two usage scenarios: (i) Unbound bokeh generation, where a user can directly sample an image at the desired blur level (see supplementary Fig. 12 for examples). (ii) Grounded bokeh generation, where a pivot image is chosen to anchor the scene whose bokeh the user wants to adjust. Since we randomly pick a pivot in each contrastive batch during training, the model learns to either add or remove blur relative to that pivot. Consequently, at inference time we can choose any pivot image with a given bokeh level and request a new target level. Figure 7 demonstrates how using either a sharper or blurrier pivot yields results matching the requested bokeh parameter without changing contents.
Real image editing.
We also support editing real photographs by first inverting them via a diffusion inversion procedure, then simultaneously denoising two latents: one to reconstruct the original image (used as pivot) and one to apply a new bokeh setting. Lower classifier-free guidance scale helps preserve details, while increasing the weight of the bokeh conditioning ensures a strong defocus effect. Figure 7 shows how our approach natively adds or reduces blur in a real photo with minimal artifacts.
4.3 Ablation Study
Figure 8: Consistency method ablation. Best viewed zoomed in.
We investigate two main aspects of our pipeline: training data composition and the method for imposing consistency.
Training data.
We train on three different settings: (i) only in-the-wild (ITW) data, (ii) only synthetic (SYN) blur augmentations, and (iii) a mixed dataset combining both. As shown in Table 2, the mixed setting achieves the best overall accuracy while also improving consistency compared to training solely on SYN data. Training on ITW data alone results in high consistency but lower accuracy. We hypothesize that the diversity and potential noise in real-world images make it harder for the model to learn a clear correlation between the conditioning signal and the depth-of-field effect. As a result, the model learns to apply only a limited bokeh generation range. Conversely, training exclusively on SYN data allows for fast convergence but degrades generation quality and generalization.
Imposing scene consistency.
We compare our complete method to the baseline conditioning method (Bokeh Cross-Attn) and a naive cross-image attention approach applied at inference (Cross-Img Attn). The naive method often forces the scene structure to remain consistent across blur conditions in comparison to baseline; however, the intended defocus variations become limited. In practice, this appears as an extremely low LPIPS (i.e., high structural consistency but also reduced bokeh range). In contrast, our grounded self-attention method, trained on paired images of differing blur, preserves both strong blur control and scene coherence, evidenced by higher correlation with the reference blur levels and closer LPIPS in Table 2. Despite this, applying cross-image attention still results in an accuracy increase (GPT-4o) relative to the baseline. We hypothesize this might derive from evaluation tendencies of GPT-4o on scene-consistent vs. inconsistent image batches. Figure 8 provides a visual comparison, demonstrating how our approach effectively maintains scene structure while still allowing for flexible depth-of-field manipulation.
5 Limitations and Discussion
Our approach relies on in-the-wild data enriched with inferred camera parameters, which inevitably introduces approximation error. For instance, our pipeline uses metric depth estimation and an empirical focus-distance estimate to decide the effective depth-of-field, yet these are derived from potentially noisy predictions. Another limitation arises from the consistency–bokeh tradeoff: to achieve highly consistent scene semantics while drastically altering the defocus parameter, the model inevitably faces tensions between rigid structural grounding and flexible blur control. Although our “grounded self-attention” helps, extreme blur changes may still introduce subtle shifts in subject contours.
Future work could expand the scope of our work, learning more flexible focal-plane controls. Likewise, we envision incorporating other photographic settings (e.g., exposure). Finally, while we primarily focus on text-to-image generation, adapting our method for image-to-image diffusion could enable stronger real-image editing use cases.
6 Conclusion
In this paper, we introduced Bokeh Diffusion, a text-to-image framework that enables precise defocus blur control while maintaining scene consistency. We integrate real in-the-wild images with synthetic blur augmentations to balance realism and robustness, while our grounded self-attention mechanism enforces structural consistency across different blur levels. Experiments show that our method surpasses baselines in both accuracy and scene consistency, offering flexible, lens-like bokeh adjustments. We hope this work inspires further research into fine-grained photographic controls in generative models.
Supplementary Material
Appendix A In-the-wild Image Augmentation
A.1 Condition Alignment
As previously discussed, real-lens defocus can be estimated via the circle-of-confusion diameter $d$ from the thin-lens model, while the synthetic blur rendering uses a scalar $K$ as the radius per unit disparity. Converting $d$ to a radius introduces a factor of $\tfrac{1}{2}$, so in pixels we have:
[…]
where $f$ and $N$ are the focal length and aperture from EXIF, $S_{1}$ is the focus distance, and pixel_ratio accounts for the camera sensor size and final image resolution.
Foreground reference and focus distance.
Since EXIF data rarely includes an explicit $S_{1}$, we approximate it by taking the median depth of the nearest portion of the foreground region. When the foreground depth range is small, we treat the entire foreground at a single $S_{1}$ to ensure a fully sharp subject. Consequently, both in-the-wild lens blur data and our synthetic disparity-driven defocus share a unified $K$ scale, enabling reliable depth-of-field conditioning across diverse real and synthetic images.
A.2 Dataset Visualization
Figure 9 illustrates the alignment between the synthetic blur augmentations and the in-the-wild defocus patterns. The figure visually compares examples from both data sources, demonstrating that the synthetic approach closely replicates the depth-of-field characteristics observed in real photographs. This alignment validates our method for unifying the blur parameter $K$ across diverse data, ensuring that the model learns consistent depth-of-field control.
Figure 9: Dataset visualization.
Appendix B Image Color Transfer
As mentioned in Section 3.4, we optionally align the color distribution of a target bokeh image with that of a reference image using $L^{*}a^{*}b^{*}$-based statistics matching. Figure 10 illustrates the effect of applying or omitting this color transfer step, demonstrating how it helps mitigate light and color biases introduced by different bokeh settings if desired.
Figure 10: Post-processing method to mitigate coloring and lighting biases sometimes introduced by different bokeh conditions.
Appendix C Additional Visual Results
Figure 11: Additional visual results for scene-consistent bokeh-conditioned generation.
Figure 12: Visual results for unbound bokeh-conditioned generation.
Figure 11 shows further scene-consistent sharp/bokeh generated pairs from our model. In Figure 12, we present visual results for unbound bokeh-conditioned generation.